# Pseudo-Stereo Inputs: A Solution to the Occlusion Challenge in Self-Supervised Stereo Matching

### v1

@@ Abstract
Self-supervised stereo matching holds great promise for application and research due to its independence from expensive labeled data. However, direct self-supervised stereo matching paradigms based on photometric loss functions have consistently struggled with performance issues due to the occlusion challenge. The crux of the occlusion challenge lies in the fact that the positions of occluded pixels consistently align with the epipolar search direction defined by the input stereo images, leading to persistent information loss and erroneous feedback at fixed locations during self-supervised training. In this work, we propose a simple yet highly effective pseudo-stereo inputs strategy to address the core occlusion challenge. This strategy decouples the input and feedback images, compelling the network to probabilistically sample information from both sides of the occluding objects. As a result, the persistent lack of information in the aforementioned fixed occlusion areas is mitigated. Building upon this, we further address feedback conflicts and overfitting issues arising from the strategy. By integrating these components, our method achieves stable and significant performance improvements compared to existing methods. Quantitative experiments are conducted to evaluate the performance. Qualitative experiments further demonstrate accurate disparity inference even at occluded regions. These results demonstrate a significant advancement over previous methods in the field of direct self-supervised stereo matching based on photometric loss. The proposed pseudo-stereo inputs strategy, due to its simplicity and effectiveness, has the potential to serve as a new paradigm for direct self-supervised stereo matching. Code is available at https://github.com/qrzyang/Pseudo-Stereo.
Index Terms:
Stereo matching, self-supervised learning, occlusion challenge
I Introduction
@@
Stereo matching, a cost-effective method for obtaining reliable 3D information, has been a longstanding focus of research in computer vision [1]. In recent years, learning-based stereo matching methods have emerged, demonstrating superior performance in terms of accuracy and efficiency [2, 3, 4]. Nevertheless, the dependence on extensive and often difficult-to-acquire annotated datasets for training poses a substantial barrier to wider adoption.
Figure 1: A comparison of various training methods in stereo matching, including a supervised method (A), the existing self-supervised paradigm (B), and our proposed self-supervised paradigm (C). (A) underscores the limitation of supervised learning, which relies heavily on accurate ground truth data. (B) illustrates that existing self-supervised methods fail to address the occlusion challenge. Our method (C) demonstrates significantly higher accuracy in the comparison within the marked boxes, showcasing both its independence from ground truth and its ability to solve the occlusion challenge.
Self-supervised stereo matching methods have emerged as a promising approach to address this challenge, attracting growing interest and research efforts. By exploiting the principle of photometric consistency across different viewpoints for network training [5, 6, 7, 8], this method eliminates the need for ground truth data. This crucial attribute notably improves the practicality and extends the scope of application for learning-based stereo matching methods. However, the photometric consistency assumption is not always valid. Typically, in occluded regions where there are no corresponding pixels from another viewpoint, the network tends to predict disparities that point towards the most similar pixels. This results in noticeable errors in disparity prediction. There has been substantial effort dedicated to addressing this issue. For example, [6] attempted to incorporate the left-right consistency assumption into the loss function, penalizing discrepancies in left and right disparity predictions. [8] discarded the losses corresponding to pixels with inconsistent left-right disparity during training, which is also a common strategy used in self-supervised optical flow estimation tasks facing similar challenges [9, 10]. Nevertheless, despite these efforts, errors in predicting occluded regions persist and notably limit the accuracy of their predictions. The main cause of this phenomenon is the absence of information in occluded regions. Neither left-right consistency nor smoothness loss can compensate for this missing information. [11] attempted to supplement the missing information by leveraging planar relationships of pixels, observing moderate success. However, their method is overly reliant on the results of edge detection, and forcing non-edge pixels into a planar structure is not consistently accurate. Consequently, existing direct self-supervised stereo matching methods that rely on photometric consistency continue to be significantly hindered by occlusion issues, leading to limited performance.
It is worth noting that self-supervised depth estimation tasks, which also utilize stereo image pairs for feedback, encounter similar challenges. However, as depth estimation only uses a single image as network input, a simple flip augmentation strategy can largely mitigate this problem [12]. Specifically, horizontal flipping can prevent occlusion regions from consistently appearing on a single side of the occluding object. This allows monocular networks to have a certain probability of obtaining information from both sides of the occluding object. However, stereo matching networks require a stereo image pair as input, with the relative position of the left and right images being fixed. This makes it impossible to directly implement this strategy. This limitation was also mentioned and analyzed in detail in [13]. This notably allows self-supervised monocular depth estimation methods to achieve significantly better performance in occluded regions compared to self-supervised stereo matching methods. This observation has led to the development of several methods that leverage depth estimation results to train stereo matching networks. [14] leverages single images and depth maps predicted by depth network to generate a stereo image dataset with disparity labels. [13] trains a monocular completion network using the same training method employed for self-supervised depth estimation. A Consensus mechanism filters and refines the predictions from this network, which then serve as ground truth labels for the supervised training of a stereo matching network. These methods outperform previous direct self-supervised approaches based on the photometric loss. However, they also significantly increase the complexity of training and reliance on monocular depth estimation, which diminishes practical applicability.
This work presents a straightforward strategy to overcome the challenge of consistently lacking information in occluded regions for self-supervised stereo matching. Specifically, we recognize the fundamental difference between stereo and monocular networks in this issue: monocular inputs can leverage feedback images independent of the input itself, while in stereo networks, the relative positions between input and feedback image pairs are fixed. This leads to consistently unchanged locations of missing information in self-supervised stereo matching. Recognizing this fundamental distinction, it becomes clear that the core challenge lies in making the input and feedback images independent for stereo networks. Given that the feedback images must be reliable and thus should not be arbitrarily adjusted, we focus our attention on the input images. Based on this principle, we introduce the pseudo-stereo inputs strategy. This approach substitutes the original inputs with generated inputs that contain a pseudo-image $\widetilde{I}^{R}$, created using the estimated disparity map output by the training network at the current time step. In contrast to methods based on dataset generation [14, 13], our approach does not rely on any previously trained networks. Pseudo-image generation solely relies on the current training network, with image accuracy improving iteratively throughout the training process. Further details are presented and discussed in Section III-A. After sequentially addressing the feedback conflicts and overfitting issues in Section III-B, this method demonstrates a stable and significant performance improvement compared to previous paradigms. By employing this approach, our self-supervised pipeline achieves competitive performance with the earlier representative supervised method DispNetC [2], without utilizing any previously trained models, post-processing techniques, or knowledge distillation strategies. This has significantly advanced the benchmarks for existing direct self-supervised stereo matching methods. In summary, the main contributions of this work are as follows.
@@
    A pseudo-stereo inputs strategy is introduced to address the persistent occlusion challenge that limits self-supervised stereo matching development.
@@
    Several other measures are proposed to further address the feedback conflicts and overfitting problem, resulting in a stable performance improvement.
@@
    This work provides a straightforward implementation that demonstrably improves the performance of direct self-supervised stereo matching methods, establishing it as a new potential paradigm.
II Related Work
II-A Learning-Based Stereo Matching Networks
Deep learning-based stereo matching methods have demonstrated superior accuracy compared to traditional approaches, leading to a surge in research interest in recent years [15]. A diverse range of methods and network architectures with unique capabilities have emerged. For example, DispNetC [2], leveraging correlation layers, boasts high matching speed. PSMNet [3], employing a 4D cost volume, achieves superior accuracy. Recent advancements have further boosted performance through different approaches. Approaches like Los [4], which emphasizes local structure information, and Temporally Consistent Stereo Matching [16], which incorporates temporal consistency, have achieved remarkable results, exhibiting both higher accuracy and faster processing speeds. Among the aforementioned methods, PSMNet, a well-established and widely-used network architecture, is frequently employed as a base network for self-supervised methods to validate their effectiveness [14, 13, 17]. Therefore, we also adopt PSMNet as our backbone in this study to facilitate comparisons.
II-B Self-Supervised Stereo Matching
II-B1 Direct Methods based on Photometric Loss
Direct methods based on photometric loss have been widely adopted in self-supervised stereo matching [5, 7, 8, 11, 18], self-supervised depth estimation [19, 20, 12, 21], and self-supervised optical flow estimation [9, 22, 10]. These approaches rely on the principle that images captured from different viewpoints should exhibit consistent brightness and color. By warping images from secondary viewpoints to align with the primary viewpoint, the methods calculate losses based on discrepancies in both photometric properties and structural details. Nevertheless, such approaches have consistently suffered from precision limitations due to factors like occlusion, which violate the assumptions of photometric consistency. In self-supervised stereo matching, various methods have been attempted to address this issue. Typical approaches include estimating and blocking the propagation of loss at occluded positions [8], penalizing inconsistencies between left and right disparity predictions [6], and exploiting adjacent information within a single image to infer disparities in occluded regions [11]. Despite these efforts, their matching performance remains constrained by occlusion problems and struggles to improve significantly.
Figure 2: Overview of the pseudo-stereo inputs strategy. The core elements encompass the pseudo-image generation method, the pseudo-stereo inputs training strategy using pseudo-images, along with the occlusion estimation and gradient detaching that aimed at resolving feedback conflicts, as well as additional methods for addressing overfitting issues (not illustrated here).
II-B2 Methods based on Dataset Generation
Training networks on synthetic datasets, which are readily available with labels, is a common practice [2]. However, due to the discrepancies between synthetic and real-world datasets, models often struggle to achieve high performance when deployed in real-world scenarios. Consequently, methods based on dataset generation have emerged that aim to generate labeled datasets from real images. [14] employs a previously trained depth estimation model to generate depth maps from single images. By combining these depth maps with the original images, they create synthetic stereo pairs and corresponding disparity labels. [13] trained a monocular completion network using the self-supervised training methods based on photometric loss. This network takes sparse disparity maps obtained through traditional methods and the original images as input, outputting dense disparity maps. The trained network can then be used to generate labeled datasets. The generated disparity maps are further refined using a consensus mechanism to remove unreliable disparity values and this distillation process significantly improves the quality of the labels, leading to superior performance for networks trained using the generated datasets. Additionally, [17] also achieved good matching accuracy by generating datasets using NeRF (Neural Radiance Fields). However, as NeRF requires numerous images of a specific scene with camera pose information as inputs, this method cannot be directly applied to common datasets like KITTI, unlike the methods previously discussed.
Overall, in self-supervised stereo matching, methods based on dataset generation have significantly outperformed direct methods based on photometric loss. The pseudo-stereo inputs strategy proposed in this paper changes this landscape. By applying this strategy, we can achieve competitive performance with a substantially lower training complexity compared to dataset generation-based approaches.
III Method
@@
The overall framework of our method is illustrated in Figure 2. Section III-A details the core idea in our approach: a pseudo-stereo inputs strategy that decouples input images from feedback images. Section III-B further discusses the additional challenges encountered during implementation and proposes solutions, including addressing feedback conflicts and overfitting problems. The final loss functions are also presented in this section.
III-A Pseudo-Stereo Inputs Strategy
@@
This section elucidates the principles and implementation of our pseudo-stereo inputs strategy, which addresses the occlusion challenge by decoupling the input image pairs from the feedback image pairs in self-supervised training. The core method for this decoupling is to utilize image pairs containing a generated pseudo-image as inputs instead of original image pairs. In contrast, the feedback calculation relies solely on the original images to ensure the reliability and accuracy of the feedback. This decoupling removes the fixed relative positional constraints on the input and feedback images, allowing them to flexibly select positions where occlusions occur, thereby addressing the issue of continuous information loss due to fixed occlusions. The following subsections provide a detailed discussion.
@@
III-A1 Pseudo-Stereo Inputs
@@
Figure 3 visually illustrates the distinction between this strategy and existing self-supervised paradigms. In this strategy, the left view remains a real image from the dataset, while the right view is replaced with a generated pseudo-image. Specifically, existing methods use a stereo image pair $I^{L}$,$I^{R}$ from the dataset as network inputs and output a disparity map $D^{L}$ aligned with $I^{L}$. Where $I$ and $D$ denote color images and disparity maps, respectively. Superscripts $L$ and $R$ indicate correspondence to the original left and right viewpoints. In contrast, when using pseudo-inputs, an estimated disparity $D^{R}$ and $I^{R}$ are used to generate a pseudo-image $\widetilde{I}^{R}$. At this point, the network inputs become $I^{R}$ as the left view and $\widetilde{I}^{R}$ as the right view, producing a dense disparity map $D^{R}$ aligned with $I^{R}$. As mentioned previously, the images used for feedback should be the original ones to ensure reliability. Therefore, the original image $I^{L}$ is warped using $D^{R}$ to obtain $I_{Warp}^{L}$, which is aligned with $I^{R}$. $I^{R}$ and $I_{Warp}^{L}$ are then employed for loss calculation. When $I^{L}$ is as the left input, the occluded areas appear to the left of the occluders; conversely, when $I^{R}$ is as the left input, the occluded areas are on the right side of the occluders. This strategy ensures that the network can consistently have a probability to capture information from both sides of the occluders, thus resolving the problem where previous methods invariably lacked information from one side of the occluders.
III-A2 Discussion on Using Pseudo-Stereo Images as Inputs
Our method differs from data generation-based methods [14, 13, 17] which use a previously trained network to generate data. Instead, the generation of pseudo-images relies on the network that is currently undergoing training and occurs throughout the entire training process. In other words, images generated in the early stages of training, which severely lack realism, are also used as inputs during the training process. This raises concerns about the network’s convergence. However, in fact, it is desirable during stereo matching network training to force the network not to overly rely on binocular information and instead seek clues within a single image. As an example, in the supervised training method [23], random patches are used to occlude the right view to accomplish this objective. Our strategy encourages the network to leverage monocular cues in the initial stages of training to enhance prediction accuracy. As accuracy increases, the precision of pseudo-image generation also improves, leading to better binocular information for the network. This results in a stable and progressively enhancing iterative process. The experimental results in Section IV-B also support this point. In these experiments, even when no real images were used for the right input (i.e., using $I^{L}$, $\widetilde{I}^{L}$ or $I^{R}$, $\widetilde{I}^{R}$ as network inputs), the network still converges successfully. This outcome is consistent with our expectations.
Figure 4: Schematic diagram of the pseudo-images generation using the rendering method, where PSMNet is the model currently under training rather than a previously well-trained one. The subscript “Flip” denotes horizontal flipping.
III-A3 Method for Generating Pseudo-Images

### v2

@@ Abstract
Self-supervised stereo matching holds great promise by eliminating the reliance on expensive ground-truth data. Its dominant paradigm, based on photometric consistency, is however fundamentally hindered by the occlusion challenge—an issue that persists regardless of network architecture. The essential insight is that for any occluders, valid feedback signals can only be derived from the unoccluded areas on one side of the occluder. Existing methods attempt to address this by focusing on the erroneous feedback from the other side, either by identifying and removing it, or by introducing additional regularities for correction on that basis. Nevertheless, these approaches have failed to provide a complete solution. This work proposes a more fundamental solution. The core idea is to transform the fixed state of one-sided valid and one-sided erroneous signals into a probabilistic acquisition of valid feedback from both sides of an occluder. This is achieved through a complete framework, centered on a pseudo-stereo inputs strategy that decouples the input and feedback, without introducing any additional constraints. Qualitative results visually demonstrate that the occlusion problem is resolved, manifested by fully symmetrical and identical performance on both flanks of occluding objects. Quantitative experiments thoroughly validate the significant performance improvements resulting from solving the occlusion challenge.
1 Introduction
@@
Stereo matching, a cost-effective method for obtaining reliable 3D information, has been a longstanding focus of research in computer vision [26]. In recent years, learning-based stereo matching methods have emerged, demonstrating superior performance in terms of accuracy and efficiency [20, 1, 18]. Nevertheless, the dominant paradigm remains fully reliant on ground-truth annotations, fundamentally limiting its potential. Self-supervised stereo matching [37, 16, 33, 31] was introduced to resolve this annotation issue and initially progressed in tandem with supervised methods, but its development soon reached a bottleneck, hindering further advancement.
The guiding principle of self-supervised stereo matching is the photometric consistency between different views [6]. This principle, however, inherently fails in regions where the consistency assumption is violated, resulting in erroneous feedback signals. Occlusion are the most critical challenge, as they produce large-scale, persistent erroneous feedback that severely degrades network performance. Consequently, a significant body of work has focused on addressing this problem [6, 16, 3, 17, 35], including related domains facing the same challenge [22, 32, 12, 10]. Their approaches can be broadly summarized into a two-pronged strategy: one is to identify occluded regions to prevent the backpropagation of erroneous gradients, and the other is to compensate for the missing information in these areas by introducing additional sources of feedback.
Information restoration for occluded areas remains a major challenge, and existing methods have yet to achieve satisfactory results. Mainstream approaches rely on providing auxiliary feedback through mechanisms like an edge-aware smoothness loss [6, 12, 31], while others attempt to exploit local disparity regularities [17, 35]. Fundamentally, these methods depend on information from the immediate neighborhood, which presumes a high degree of regularity and correlation between the occluded region and its surroundings—an assumption that is often unreliable. Some works [6, 16, 4] advocate for using a flip data augmentation to source information from symmetric regions. However, as demonstrated in Fig. 1, this strategy is ineffective. Due to the coupled nature and fixed directionality of the input and feedback signal, occlusions consistently appear on a fixed side of the occluding object in the primary view, and the flip operation fails to alter this geometric reality. In contrast, methods [36, 28] that introduce a third viewpoint—using a central primary view and selecting the minimum loss from the left and right views—have proven effective at correctly resolving occlusions. Nevertheless, these solutions are impractical for standard settings. They either necessitate specialized, perfectly symmetric trinocular camera hardware [36] or require a full 3D scene reconstruction via methods like NeRF to synthesize the novel view [28]. Consequently, such techniques cannot be directly applied for self-supervised training on conventional binocular image data.
Figure 1: An illustration of our motivation. (Top) The fixed-side occlusion problem where standard flipping is ineffective, as the occlusion side remains constant relative to the inputs. (Bottom) Our method decouples inputs and feedback using a pseudo-view (C), breaking the persistent directional error signal.
In this work, we propose a novel occlusion handling method that acquires reliable feedback, instead of relying on simple inference from local neighbors, and also without introducing any additional constraints. The core idea, illustrated in Fig. 1, is to decouple the network input and its photometric feedback signal by using a pseudo third-view input. This principle effectively reframes the problem from one of static, one-sided occlusions to one of dynamic, two-sided occlusions with equal probability. The execution of this idea is non-trivial, and we identified and solved three core challenges. First, how to generate the pseudo third-view input with minimal cost and using it to achieve the aforementioned transformation to symmetric occlusions. Second, given that symmetric occlusions imply that each side of an occluder receives correct and incorrect feedback with equal probability, how to guide the network to converge to the correct solution. Third, how to ensure proper network convergence and prevent overfitting, given the distinguishable nature of the pseudo-input compared to the real input.
Our experiments validate that the proposed full framework reliably resolves the occlusion problem, manifested by the symmetric and accurate estimates on both sides of occluders shown in Fig. 6, and the significant performance gains reported in Tab. 1. Furthermore, subsequent experiments (Fig. 7 and Tab. 2) confirm that occlusion is a core bottleneck in self-supervised stereo matching, independent of network architecture or dataset. We demonstrate that our proposed method successfully overcomes this fundamental bottleneck across diverse network architectures and datasets.
In summary, the main contributions of this work are as follows.
@@
    We propose a novel occlusion handling strategy that uses pseudo-stereo inputs to decouple network input and feedback. This fundamentally reframes the learning problem in occluded regions.
@@
    We introduce a complete and robust framework to realize this idea, ensuring convergence under feedback conflicts and preventing overfitting to synthetic artifacts.
@@
    Extensive experiments show our method resolves the occlusion bottleneck to achieve significant performance improvements, generalizing effectively across various network backbones and datasets.
2 Related Work
Figure 2: An overview of our proposed pipeline. We leverage a pseudo-input to decouple network inputs and feedback. This enables a symmetric training paradigm where occlusions are probabilistically relocated, providing valid feedback for both sides of occluders. An occlusion estimation and gradient detaching mechanism is employed to resolve the inherent feedback conflicts.
2.1 Learning-Based Stereo Matching Backbone
The rapid advancement of deep learning has given rise to powerful stereo matching architectures [8, 29]. Representative works like PSMNet [1] introduced 3D convolutions for cost aggregation, while more recent methods like GMStereo [34] incorporate transformers [30] for more capable feature matching. Recently, approaches such as MonSter [2] have shown success by integrating features from pre-trained monocular depth estimators. Despite their increasingly powerful feature representation and matching capabilities, these models are fundamentally designed for a supervised paradigm and require accurate ground-truth disparities. When applied to self-supervised learning, where the supervision signal can be ambiguous or erroneous, their advanced feature induction capabilities, while still beneficial, do not address the core bottleneck. Although some works [16, 31, 3] have proposed architectural modifications specifically for self-supervised occlusion handling, their visual results reveal that the underlying problem is not significantly mitigated. This suggests that the solution to the occlusion problem lies primarily within the loss formulation rather than the network architecture itself.
2.2 Self-Supervised Stereo and Occlusion Handling
Self-supervised stereo methods are built upon the principle of photometric consistency between views [37, 33, 31, 35, 4]. Consequently, a large body of work has focused on strategies to address this principle’s primary weakness: the erroneous feedback generated by occlusions. A common consensus is to identify occluded regions via a mask and exclude them from the photometric loss, typically through pixel-wise multiplication [22, 32, 33, 12, 31]. Method [12] compared various occlusion inference methods [32, 22, 11], demonstrating that all benefit performance, and crucially established that the gradient stop for occlusion mask is an important strategy.
With occlusions masked, however, an information void is created. Various methods have been proposed to compensate for this. Without introducing external constraints, this compensation is typically provided by a standard smoothness loss [31, 22, 32, 33, 12]. [12] demonstrated that different smoothness edge-weights can significantly impact performance. Nevertheless, since smoothness is not always a reliable or correct convergence target, this approach remains unsatisfactory. Some methods [6, 16, 4] employ flip strategy and regard left-right consistency directly as a loss term, attempting to use it as a key solution. However, this does not fundamentally mitigate the occlusion problem, as an incorrect disparity can still satisfy the consistency. A more targeted line of work [17, 35] have designed particular losses to extract more reliable information from local neighborhoods. Among these, [17] focuses on interpolating from background neighbors rather than foreground ones, while [35] encourages occluded regions and their neighborhoods to conform to a 3D geometry inferred from monocular cues. These methods are more tailored for occlusion handling and have achieved performance gains, but their solutions are not robust as they rely on the strong assumption that neighboring information is sufficiently correlated with the occluded content.
Distinct from the aforementioned methods, some works [36, 28] acquire reliable, non-local feedback by leveraging additional views. Specifically, [36] uses a fixed, symmetric multi-view camera hardware where a region occluded from one direction is visible from another. [28] achieves the same goal by rendering a third-view image using Neural Radiance Fields [24] (NeRF). These two approaches provide a conceptually correct pathway for occlusion handling: obtaining unoccluded information from an auxiliary perspective. However, both methods either depend on specialized hardware or require additional data for scene reconstruction, imposing external constraints that prevent their direct application to conventional binocular datasets. This is inconsistent with the goal of self-supervised learning, which aims for direct training on any target scene.
In this paper, we propose a self-supervised method that can acquire the same reliable feedback as these multi-view approaches without depending on any such external constraints.
3 Method
@@
The motivation and the overall framework of our method are illustrated in Fig. 1 and Fig. 2, respectively. Section 3.1 details the core idea in our approach: a pseudo-stereo inputs strategy that decouples input images from feedback images. Section 3.2 further discusses the practical challenges of this design and proposes solutions, including addressing feedback conflicts and overfitting problems.
3.1 Pseudo-Stereo Inputs Strategy
@@
This section elucidates the principles and implementation of our pseudo-stereo inputs strategy, which addresses the occlusion challenge by decoupling the input image pairs from the feedback image pairs in self-supervised training. The core method for this decoupling is to utilize image pairs containing a generated pseudo-image as inputs instead of original image pairs. In contrast, the feedback calculation relies solely on the original images to ensure the reliability and accuracy of the feedback. This decoupling removes the fixed relative positional constraints on the input and feedback images, allowing them to flexibly select positions where occlusions occur, thereby addressing the issue of continuous information loss due to fixed occlusions. An intuitive illustration of the principle is shown in Fig. 1.
@@
3.1.1 Pseudo-Stereo Inputs
@@
Figure 3 visually illustrates the distinction between this strategy and existing self-supervised paradigms. In this strategy, the left view remains a real image from the dataset, while the right view is replaced with a generated pseudo-image. Specifically, existing methods use a stereo image pair $I^{L}$$I^{R}$ from the dataset as network inputs and output a disparity map $D^{L}$ aligned with $I^{L}$. Where $I$ and $D$ denote color images and disparity maps, respectively. Superscripts $L$ and $R$ indicate correspondence to the original left and right viewpoints. In contrast, when using pseudo-inputs, an estimated disparity $D^{R}$ and $I^{R}$ are used to generate a pseudo-image $\widetilde{I}^{R}$. At this point, the network inputs become $I^{R}$ as the left view and $\widetilde{I}^{R}$ as the right view, producing a dense disparity map $D^{R}$ aligned with $I^{R}$. The images used for feedback should be the original ones to ensure reliability. Therefore, the original image $I^{L}$ is warped using $D^{R}$ to obtain $I_{Warp}^{L}$, which is aligned with $I^{R}$. $I^{R}$ and $I_{Warp}^{L}$ are then employed for loss calculation. When $I^{L}$ is as the left input, the occluded areas appear to the left of the occluders; conversely, when $I^{R}$ is as the left input, the occluded areas are on the right side of the occluders. This strategy ensures that the network can consistently have a probability to capture information from both sides of the occluders, thus resolving the problem where previous methods invariably lacked information from one side of the occluders.
3.1.2 Method for Generating Pseudo-Images